OpenStudy (anonymous):

Find the set of δ values that satisfy the formal definition of lim(x→4)⁡〖11/(g(x))〗=2.2 when given the value ε = 0.5, showing all work. Knowing that g(x)= x^2-x-7

OpenStudy (anonymous):

\[\lim_{x\to4}\frac{11}{x^2-x-7}=2.2~~?\]

OpenStudy (anonymous):

Nevermind, I solved it. It is using a delta epsilon proof (calc III) so it is intensive. Thanks Anyway

OpenStudy (anonymous):

To find the proper \(\delta\), assume \(\left|\dfrac{11}{x^2-x-7}-\dfrac{11}{5}\right|<\epsilon\). Then, \[\begin{align*}\left|\dfrac{11}{x^2-x-7}-\dfrac{11}{5}\right|&=11\left|\dfrac{5-(x^2-x-7)}{x^2-x-7}\right|\\ &=11\left|\frac{x^2-x-12}{x^2-x-7}\right|\\ &=11\left|\frac{(x-4)(x+3)}{x^2-x-7}\right|\\ &=11\left|x-4\right|\left|x+3\right|\left|\frac{1}{x^2-x-7}\right| \end{align*}\] Agree to set \(\delta\le1\), then if \(0<|x-4|<\delta\), you have that \[\begin{align*} |x-4|&<1\\ -1<x-4&<1\\ 6<x+3&<8\\ |x+3|&<8 \end{align*}\] So, \[11\left|x-4\right|\left|x+3\right|\left|\frac{1}{x^2-x-7}\right|<88\left|x-4\right|\left|\frac{1}{x^2-x-7}\right|\] For the quadratic factor, I think completing the square would be your best bet: \[x^2-x-7=x^2-x+\frac{1}{4}+\frac{27}{4}=\left(x-\frac{1}{2}\right)^2+\frac{27}{4}\] So if \(\delta\le1\), you get \[\begin{align*} |x-4|&<1\\ -1<x-4&<1\\ \frac{5}{2}<x-\frac{1}{2}&<\frac{9}{2}\\ \left|x-\frac{1}{2}\right|&<\frac{9}{2}\\ \left(x-\frac{1}{2}\right)^2&<\frac{81}{4}\\ \left(x-\frac{1}{2}\right)^2+\frac{27}{4}&<\frac{108}{4}=27\\ \frac{1}{\left(x-\frac{1}{2}\right)^2+\frac{27}{4}}&>\frac{1}{27} \end{align*}\] So, \[88\left|x-4\right|\left(\frac{1}{27}\right)<88\left|x-4\right|\left|\frac{1}{x^2-x-7}\right|<\epsilon\] or simply \[\frac{88}{27}\left|x-4\right|<\epsilon\\ |x-4|<\frac{27\epsilon}{88}\] Which would mean you would have to set \(\delta=\min\left\{1,\dfrac{27\epsilon}{88}\right\}\). So if \(\epsilon=\dfrac{1}{2}\) is given, you would pick \(\delta=\min\left\{1,\dfrac{44}{27}\right\}=1\). Now, I get the feeling I might be missing something in the proof, but I'm comfortable with it overall.

OpenStudy (anonymous):

... Just noticed that you have your answer... Oh well :)